# Non-Hermitian higher-order topological insulators enabled by altermagnet engineering

### v1

@@ Abstract
We show that proximitizing an altermagnet to a non-Hermitian topological insulator provides a powerful mechanism for engineering non-Hermitian higher-order topological phases. The altermagnetic order opens a gap at the topological edge states and drives a topological phase transition from a first-order to a second-order topological phase. When combined with nonreciprocal hopping, the system exhibits both the non-Hermitian skin effect and a hybrid skin–topological effect, whereby first-order edge states and second-order corner states accumulate at selected corners of the lattice. We demonstrate that the spectral winding number of the edge states under cylindrical geometry dictates this corner localization and can be reversed by tuning the altermagnetic order. Consequently, both edge and corner states become directionally controllable. Our results establish altermagnets as a versatile platform for realizing and tuning skin-topological phenomena in non-Hermitian higher-order topological systems.
@@ I Introduction
Altermagnetism has recently emerged as a distinct class of collinear magnetic order that bridges the physics of ferromagnets and antiferromagnets [1, 2, 3, 4, 5, 6, 7, 8]. It exhibits large momentum-dependent spin polarization while maintaining zero net magnetization, thus realizing time-reversal–odd spin splitting with compensated magnetic moments [9, 10, 11]. This unique symmetry realization enriches the taxonomy of magnetic phases and provides an intrinsic route to dissipationless spin control without stray fields. Its symmetry-filtered spin texture enables precise band-structure engineering, as verified theoretically and experimentally in RuO2 [12, 13, 14], MnTe [15, 16, 17], CrSb [18, 19, 20], Mn5Si3 [21, 22], and KV2Se2[23]. Due to its time-reversal breaking band structure and alternating momentum-dependent spin splitting, altermagnetism gives rise to a variety of novel effects and phases, including the crystal Hall effect [24, 25, 26, 27], giant tunneling magnetoresistance [28, 29], spin–momentum interactions [30, 31, 32], and topological phases [33, 34, 35, 36, 37].
In recent years, the study of topological phases has been extended to non-Hermitian systems, attracting substantial research interest [38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58]. Non-Hermitian systems exhibit a variety of striking phenomena that have no counterpart in Hermitian physics [59, 60, 61, 62, 63]. Among the most prominent are exceptional points [64, 65], where both eigenvalues and eigenvectors coalesce, and the non-Hermitian skin effect (NHSE) [42, 66, 67, 68, 69, 70, 71, 72, 73, 74], which drives an extensive number of bulk states to accumulate at system boundaries under open boundary conditions(OBC). By reshaping the topology of complex energy spectra, non-Hermiticity enables unconventional topological phenomena, especially in higher dimensions. In particular, the hybridization of topological edge states with the NHSE gives rise to the skin–topological effect, where topologically protected edge states localize directionally at lower-dimensional boundaries [75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85]. These phenomena highlight the importance of understanding how non-Hermiticity interacts with intrinsic topological boundary states. Exploring this interplay in altermagnetic systems is especially compelling: while recent works have examined exceptional-point physics in non-Hermitian altermagnets [86, 87], the NHSE and its coupling to higher-order topology in altermagnetic platforms remain largely unexplored.
@@ II Switchable Skin-topological effect
Here, $H_{\text{TI}}(\mathbf{k})=(m_{0}-t_{x}\cos k_{x}-t_{y}\cos k_{y})\sigma_{x}+A_{x}\sin k_{x}\sigma_{z}s_{x}+A_{y}\sin k_{y}\sigma_{z}s_{y}$ describes a Hermitian topological insulator that preserves time-reversal symmetry and hosts helical edge states, relevant to materials like HgTe and WTe2 [88, 89, 90]. The non-Hermiticity is introduced by $H_{\text{NH}}(\mathbf{k})=(-i\delta_{x}\sin k_{x}-i\delta_{y}\sin k_{y})\sigma_{x}$, which effectively models a nonreciprocal hopping arising from an imaginary gauge field. The proximity-induced altermagnetic order is given by $H_{\text{AM}}(\mathbf{k})=J(\cos k_{x}-\cos k_{y})\sigma_{z}s_{z}$, characterizing a $d$-wave spin splitting where $J$ denotes the strength of the altermagnetism. $s_{i}$ and $\sigma_{i}$ denote the Pauli matrices acting on spin and orbital degrees of freedom, respectively.
@@ IV Conclusion
In this work, we have investigated the a non-Hermitian second-order topological insulator realized by placing a non-Hermitian topological insulator in proximity to an altermagnet. The nonreciprocal hopping drives the bulk states to accumulate at the lattice corners, generating a pronounced non-Hermitian skin effect. Meanwhile, the altermagnetic order gaps the topological edge states and induces a transition from a first-order to a second-order topological phase. The interplay between altermagnetism and non-Hermiticity produces a skin–topological effect, in which the first-order edge states are driven to the corners of the lattice. This effect is accurately captured by the spectral winding number of the edge states under cylindrical geometry. Crucially, the proximitized altermagnetic order reshapes the real-part edge dispersion, enabling a sign reversal of the spectral winding number and thus a controllable inversion of edge and corner localization, without altering the direction of nonreciprocal hopping. As a consequence, the second-order corner states inherit the skin–topological localization and preferentially accumulate at the same corners as the edge states. Our results reveal the rich interplay between altermagnetism and non-Hermiticity, demonstrating that non-Hermitian altermagnets offer a versatile and tunable platform for manipulating topological edge and corner states.
@@
V Acknowledgments
This work was supported by Natural Science Foundation of Jiangsu Province (Grant No. BK20231320).

### v2

@@ Abstract
We show that proximity to an altermagnet provides an efficient route to engineering non-Hermitian higher-order topological phases. The proximity-induced altermagnetic order gaps the edge states of a topological insulator, thereby driving a transition from a first-order to a second-order topological phase. When combined with nonreciprocal hopping, the system exhibits both the non-Hermitian skin effect and a hybrid skin–topological effect, whereby first-order edge states and second-order corner states accumulate at selected corners of the lattice. We demonstrate that the spectral winding number of the edge states under cylindrical geometry dictates this corner localization and can be reversed by tuning the altermagnetic order. Consequently, both edge and corner states become directionally controllable. Our results establish altermagnets as a versatile platform for realizing and tuning skin-topological phenomena in non-Hermitian higher-order topological systems.
@@ I Introduction
Altermagnetism has recently emerged as a distinct class of collinear magnetic order that bridges the physics of ferromagnets and antiferromagnets [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. It exhibits large momentum-dependent spin polarization while maintaining zero net magnetization, thus realizing time-reversal–odd spin splitting with compensated magnetic moments [12, 15]. This unique symmetry realization enriches the taxonomy of magnetic phases and provides an intrinsic route to dissipationless spin control without stray fields. Its symmetry-filtered spin texture enables precise band-structure engineering, as verified theoretically and experimentally in MnTe [16, 17, 18], CrSb [19, 20, 21], Mn5Si3 [22, 23], RbV2Se2[15] and KV2Se2[24, 25]. Due to its time-reversal breaking band structure and alternating momentum-dependent spin splitting, altermagnetism gives rise to a variety of novel effects and phases, including the crystal Hall effect [26, 27, 28, 29], giant tunneling magnetoresistance [30, 31], spin–momentum interactions [32, 33, 34], unique proximity effect [35, 36] and topological phases [37, 38, 39, 40, 41].
In recent years, the study of topological phases has been extended to non-Hermitian systems, attracting substantial research interest [42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63]. Non-Hermitian systems exhibit a variety of striking phenomena that have no counterpart in Hermitian physics [64, 65, 66, 67, 68]. Among the most prominent are exceptional points [69, 70], where both eigenvalues and eigenvectors coalesce, and the non-Hermitian skin effect (NHSE) [46, 71, 72, 73, 74, 75, 76, 77, 78, 79], which drives an extensive number of bulk states to accumulate at system boundaries under open boundary conditions(OBC). By reshaping the topology of complex energy spectra, non-Hermiticity enables unconventional topological phenomena, especially in higher dimensions. In particular, the hybridization of topological edge states with the NHSE gives rise to the skin–topological effect, where topologically protected edge states localize directionally at lower-dimensional boundaries [80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90]. These phenomena highlight the importance of understanding how non-Hermiticity interacts with intrinsic topological boundary states. Exploring this interplay in altermagnetic systems is especially compelling: while recent works have examined exceptional-point physics in non-Hermitian altermagnets [91, 92], the NHSE and its coupling to higher-order topology in altermagnetic platforms remain largely unexplored.
@@ II Switchable Skin-topological effect
Here, $H_{\text{TI}}(\mathbf{k})=(m_{0}-t_{x}\cos k_{x}-t_{y}\cos k_{y})\sigma_{x}+A_{x}\sin k_{x}\sigma_{z}s_{x}+A_{y}\sin k_{y}\sigma_{z}s_{y}$ describes a Hermitian topological insulator that preserves time-reversal symmetry and hosts helical edge states, relevant to materials like HgTe and WTe2 [93, 94, 95]. The non-Hermiticity is introduced by $H_{\text{NH}}(\mathbf{k})=(-i\delta_{x}\sin k_{x}-i\delta_{y}\sin k_{y})\sigma_{x}$, which effectively models a nonreciprocal hopping arising from an imaginary gauge field. The proximity-induced altermagnetic order [96, 35], supported by multiple realistic materials, is given by $H_{\text{AM}}(\mathbf{k})=J(\cos k_{x}-\cos k_{y})\sigma_{z}s_{z}$, characterizing a $d$-wave spin splitting where $J$ denotes the strength of the altermagnetism. $s_{i}$ and $\sigma_{i}$ denote the Pauli matrices acting on spin and orbital degrees of freedom, respectively.
@@ IV Conclusion
In this work, we have investigated the non-Hermitian second-order topological insulator realized by placing a non-Hermitian topological insulator in proximity to an altermagnet. The nonreciprocal hopping drives the bulk states to accumulate at the lattice corners, generating a pronounced non-Hermitian skin effect. Meanwhile, the altermagnetic order gaps the topological edge states and induces a transition from a first-order to a second-order topological phase. The interplay between altermagnetism and non-Hermiticity produces a skin–topological effect, in which the first-order edge states are driven to the corners of the lattice. This effect is accurately captured by the spectral winding number of the edge states under cylindrical geometry. Crucially, the proximitized altermagnetic order reshapes the real-part edge dispersion, enabling a sign reversal of the spectral winding number and thus a controllable inversion of edge and corner localization, without altering the direction of nonreciprocal hopping. As a consequence, the second-order corner states inherit the skin–topological localization and preferentially accumulate at the same corners as the edge states. Our results reveal the rich interplay between altermagnetism and non-Hermiticity, demonstrating that non-Hermitian altermagnets offer a versatile and tunable platform for manipulating topological edge and corner states.
@@
V Acknowledgements
This work was supported by Natural Science Foundation of Jiangsu Province (Grant No. BK20231320), National Natural Science Foundation of China (12474155) and the Zhejiang Provincial Natural Science Foundation of China (LR25A040001).